OpenStudy (abbycross167):

Write the slope-intercept (y = mx + b) form of the equation from the information given. 3. The line goes through the points (8,5) and (9, 6)

OpenStudy (abdullahm):

So we need to find the slope first \(\sf\large Slope = \frac{y_2-y_1}{x_2-x_1} = \frac{6-5}{9-8}=?\)

OpenStudy (abbycross167):

1/1?

OpenStudy (abbycross167):

@AbdullahM ?

OpenStudy (abdullahm):

correct :)

OpenStudy (abdullahm):

now we plug in a point into y = x + b So... 5 = 8 + b what is b?

OpenStudy (abbycross167):

-3?

OpenStudy (abdullahm):

yayay correct :)

OpenStudy (abdullahm):

so your full equation is y = x - 3

OpenStudy (abbycross167):

That's it? no other numbers go in the equation?

OpenStudy (abdullahm):

none other :)

OpenStudy (abbycross167):

oh wow ok thank you!! if i post a new question and show my work could you check it?

OpenStudy (abdullahm):

sure, just tag me to the new post :)